Find the equation (in terms of x) of the line through the points (-4,3) and (3,1)

Answer by stanbon(75887)

Find the equation (in terms of x) of the line through the points (-4,3) and (3,1)
------
slope = (3-1)/(-4-3) = -2/7
-----
Form:: y = mx + b
Solve for "b":
1 = (-2/7)3 + b
(7/7) = -6/7 + b
b = 13/7
--------------
Equation:
y = (-2/7)x + (13/7)
======================
Cheers,
Stan H.
======================